# Small but large: Single organic molecules as hybrid platforms for quantum technologies

### v1

@@ III.2 Molecular optomechanics
The optical and mechanical degrees of freedom are weakly interacting beyond systems at the atomic scale due to the large mechanical mass and small momentum of optical beams. In the last decade, highly cooperative interactions between those have been made possible by utilizing engineered optomechanical cavities with low optical and mechanical losses [92]. These efficient interactions enabled precise control and manipulation of mechanical motion with light, and hold promises for quantum sensing, signal transduction, and so on [93].
@@
Analogous to the conventional optomechanical setups, a single-molecule host-guest system offers optomechanical interactions thanks to the inherent couplings between their optically accessible electronic and vibrational transitions. The molecular system provides such couplings across a very wide frequency range spanning kHz to THz. Importantly, the electron-vibration coupling strengths can reach values larger than the mechanical oscillation frequencies, putting the molecular system into the ultra-strong optomechanical coupling regime at the single-photon level [46]. This nonlinear quantum optomechanical regime has not been reached in solid-state cavity optomechanical setups so far [92]. In addition, due to the extremely large quality factors of the electronic transitions in the order of $10^{7}$, these molecular systems could also ensure a path to access large quantum optomechanical cooperativities, i.e. $C_{\text{opt}}=4g^{2}/\bar{n}\kappa_{v}\gamma_{0}$ [94], where $\bar{n}$ denotes the thermal population of the vibrational mode, $\kappa_{v}$ and $\gamma_{0}$ denotes the vibrational and electronic excited state decay rates, respectively. Reaching such a high cooperativity regime in cavity optomechanical systems is challenging due to parasitic optical absorption [94], which is negligible in molecular systems. Hence, single molecules can facilitate quantum optomechanical applications such as coherent frequency conversion between photons [95] which is essential for quantum information processing [96].
In typical single-molecule host-guest systems, the efficient optomechanical interactions suffer from the fast vibrational relaxations as the quantum character of these levels is lost in the timescale of electron-vibration couplings. Therefore, it is important to control vibrational relaxation rates. Such a control can be done for acoustic phonons as in traditional cavity optomechanical setups, where techniques like impedance mismatch or phononic bandgap structuring can be used to decouple certain acoustic modes from their environment. For anthracene crystal doped with dibenzoterrylene, it has been proposed that efficient optomechanical interactions are at reach by using appropriately designed phononic structures [38]. On the other hand, it is a great challenge to extend such a control to optical phonons and guest molecular vibrations as discussed in Section II.2. Considering their significant electron-vibration coupling strengths and non-negligible thermal populations even at room temperature, engineering such long-lived states could lead to efficient optomechanical applications at elevated temperatures. This approach also offers the potential to explore previously inaccessible regimes in optomechanical and nonequilibrium condensed-matter systems.
@@
Figure 6: (a) Achieving optical manipulation of the singlet-triplet transition will provide a “switch” to control the electron spin, allowing it to be turned “on” and “off” on demand. (b) This capability combined with coherent photon scattering via the |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ to |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ ZPL transition facilitates optical interfacing of the electron spin. Illustrated is a spin-controlled photon transmission through a cavity, whose frequency is set resonant to the ZPL transition.
@@ III.3 Spin-photon interface
The spin states are long-lived and well-isolated from their environment, making them attractive candidates as quantum memory units. Additionally, the excellent properties of molecular ZPL emission offers the possibility to coherently interface spins with propagating photons. These two features combined make it possible to realize a molecular spin-photon interface, which can serve as a key component for quantum networks [97].
To realize a coherent spin-photon interface, it is necessary not only to have optically addressable spin states but also to achieve an efficient collection of lifetime-limited photons into a single optical mode. For solid-state emitters, the efficiency of detecting such lifetime-limited photons is constrained by the emitter’s internal quantum efficiency, the branching ratio of ZPL emission, and the optical dephasing rate. An established approach to mitigate these limitations is through resonant coupling with an optical cavity [98]. PAH molecules are particularly promising in this regard due to their high quantum efficiency, close to lifetime-limited ZPL emission, and relatively high branching ratio. Recent advances in cavity-based interfaces have achieved strong coupling between the ZPL transition of a single molecule and a tunable Fabry-Perot microcavity [99, 10], laying the groundwork for the realization of a molecular spin-photon interface.
@@
Figure 6(a) illustrates a possible scheme of a molecular spin-photon interface. To entangle the electron spin with photons, one would need to prepare the molecule in a superposition of |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ and a sublevel of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$. This can in principle be achieved using a narrow linewidth laser to coherently drive the transition between these two states. The |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ to |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ transition is strongly coupled to an optical microcavity and serves as an optical interface. Depending on the state of the molecule, a photon with its frequency matching the undressed cavity frequency will either reflects off from the cavity, due to the vacuum Rabi splitting [10] when the molecule is in |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, or transmits through the cavity when the molecule is in |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ (considering a symmetric mirror configuration). Thus the internal state of the molecule becomes entangled with the spatial mode of the photon.
@@
A critical step in this scheme is the optical addressing of singlet-triplet transition in a single molecule. Such transitions have been studied in experiments with cold alkaline earth atoms, e.g., strontium, and utilized for optical clocks transitions [100]. In recent years, singlet-triplet transitions have been measured on diatomic molecules in the gas phase [101, 102]. Optical excitation of singlet-triplet transitions in a single solid-state emitter has not been demonstrated. The challenges are mainly technical. First, in planar molecules without heavy metal atoms, the spin-orbit coupling is rather weak, leading to long lifetimes and thus narrow transition linewidths. Second, the exact frequencies of these transitions are difficult to predict, and the matrix introduces shifts to the transition frequency compared to that of a molecule in vacuum. In addition, considerations need to be taken in identifying molecular candidates for such an investigation. We will discuss possible molecular candidates in Section III.5.
@@ III.4 Molecular spin register
Apart from nuclear spin qubits, using nuclei with larger spin quantum numbers can extend the system to a qudit (a system with a $d$ number of levels) register. Nuclear spins with higher multiplicity offer an interesting route for quantum information processing beyond qubit-based approaches [110]. They are predicted to bring along reduced numbers of physical quantum units and gate operations for accomplishing a certain computing task. A parallel line of research on molecular qudits is being pursued in the context of molecular nanomagnets [110]. On the PAH platform, it is readily feasible to incorporate a defined number of spin-1 nuclei, e.g., 15N, through bay functionalization with imide groups [111]. Along this line, perylenediimide and terrylenediimide derivatives have already been studied in single-molecule optical spectroscopy [112, 113].
@@
Overall, the molecular framework provides the capability to integrate electron spin, nuclear spins, and optical interfaces within a single platform. In addition to the tailorability by chemical modification, a particularly appealing aspect of this system is the metastable triplet state. Once optical driving of such a transition is realized, it enables a deterministic control to switch “on” and “off” the electron spin [see red arrow in Fig. 6(a)]. This feature is particularly compelling, given that the electron spin acts as a significant source of decoherence for the nuclear spins [114]. By enabling a deterministic on-off switch of the hyperfine interaction, this system can address the trade-off between gate time and susceptibility to decoherence, further enhancing the utility of the molecular platform.
@@
III.5 Exploiting the Chemical Space
@@
We provide a glimpse into the chemical space of molecules in Fig. 7, where $26$ centrosymmetric PAH molecules containing up to $26$ carbon atoms are identified from the NIST database [115]. Their singlet-singlet (|$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$) and singlet-triplet (|$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$) vertical transition energies are calculated using ab initio methods at the level of coupled-cluster theory with single and double excitations (CCSD) [123] and plotted versus each other. Note that these energies do not include salvation effects or zero-point energy correction stemming from the vibrational degrees of freedom. The results show an approximately linear scaling behavior between the singlet-singlet and singlet-triplet transition energies (see dashed black line in Fig. 7), spanning a large range in the visible and near-infrared spectrum. This molecular space has not been studied at the single-molecule level except for perylene and pentacene. We discuss the potential of such a diverse chemical space with two possible applications.
@@
(i) The large variety of |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ transition energies of these molecules offers possibilities to engineer single photon sources with on-demand emission frequencies across the visible spectrum. Although the identified molecules do not cover the entire range, the wavelength gaps between them could be bridged using several strategies. First, chemical modifications, such as deuterium and fluorine substitution of protons, or the adaption of imide or anhydride groups, can shift the center emission wavelength by up to $\sim$50 nm [127, 112, 113]. Second, selecting different host matrices or different insertion sites in a host can affect the emission wavelength by up to $20$ nm [128]. Third, within a given host matrix, the center of inhomogeneous broadening can be tuned with the thickness of the host crystal, providing an accessible frequency range of several nanometers [129]. Fourth, the inhomogeneous broadening of ZPL frequencies can span over $\sim$1 nm, allowing spectral selection of molecules within this frequency range. Finally, fine adjusting of the ZPL emission frequency can be achieved through the Stark effect, with a tunning range of $\sim$500 GHz [12].
@@
(ii) The molecules in Fig. 7 also offers a valuable chemical space for identifying molecular candidates suitable for realizing spin-photon interfaces, where the optical addressing of the singlet-triplet transition is essential but remains an ongoing challenge [128]. Molecules suitable for addressing singlet-triplet transitions should meet two conditions regarding their transition energy: First, the energy-gap law predicts that a lower energy of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ is associated with a larger non-radiative ISC rate [79]. Therefore, molecules with a higher |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ energy are more favorable. Second, since the energy of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ is lower than that of |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, a high energy of |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ poses challenges. This makes it difficult to find suitable host materials and necessitates widely tunable laser sources in the ultraviolet, which is technically demanding. Therefore, one would aim for molecules with a high |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ energy but with a relatively small energy gap to the |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ state. These two criteria let us constraint the selection of guest molecules into the subspace illustrated by the blue-shaded area in Fig. 7.
These two examples are just a glimpse of the immense possibilities provided by the chemical space. A larger chemical space can be screened in the future using machine learning strategies in theoretical [130] or experimental settings [131] to identify new host and guest combinations with specific properties such as ZPL frequency at telecom wavelengths, long vibrational lifetimes, large electron-vibration coupling strengths or spin coherence times. Eventually, inverse molecular design techniques [126] can be employed to create new molecules with tailored photonic, vibronic, and spin properties, unleashing the full power of chemical synthesis. These approaches could drive advances beyond traditional intuitive chemical design principles and accelerate the development of molecules for quantum technologies.

### v2

@@ III.2 Molecular optomechanics
Optical and mechanical degrees of freedom typically interact weakly in systems beyond the atomic scale. This is due to the large mechanical mass, which makes mechanical systems less responsive to the small momentum of optical fields. In the last decade, highly cooperative interactions between those have been made possible by utilizing engineered optomechanical cavities with low optical and mechanical losses [102]. These efficient interactions enabled precise control and manipulation of mechanical motion with light, and hold promises for quantum sensing, signal transduction, and so on.
@@
Analogous to conventional optomechanical setups, a single-molecule host-guest system offers optomechanical interactions thanks to the inherent couplings between their optically accessible electronic and vibrational transitions. The molecular system provides such couplings across a very wide frequency range spanning kHz to THz. Importantly, the electron-vibration coupling strength ($g_{0}$) can reach values comparable to the mechanical oscillation frequencies of vibrons [44], i.e., $g_{0}\sim 0.1-0.5~{}\omega_{v}$, putting the molecular system into the ultra-strong optomechanical coupling regime at the single-photon level [103]. This nonlinear quantum optomechanical regime has not been reached in solid-state cavity optomechanical setups so far.
Together with the extremely large quality factors of the electronic transitions ($\gamma_{0}\sim$ MHz, where $\gamma_{0}$ is the spontaneous decay rate of electronic transitions), these ultra-strong coupling strengths enable the system to achieve a high quantum optomechanical cooperativity, $C_{\text{opt}}$, given by $C_{\text{opt}}=4g_{0}^{2}/\bar{n}\kappa_{v}\gamma_{0}$ [103]. Here, $\bar{n}$ represents the thermal population of the vibrational modes and $\kappa_{v}$ is the decay rate of the vibrational excited state. Although the vibrons have a short lifetime ($\kappa_{v}\sim 10~{}\text{ps}$), the combination of these parameters enables the realization of $C_{\text{opt}}>10^{5}$ over a wide range of temperatures ($\bar{n}\ll 1$). Reaching such a high cooperativity regime in cavity optomechanical systems is challenging due to parasitic optical absorption [104], which is negligible in molecular systems. Hence, single molecules could facilitate quantum optomechanical applications such as coherent frequency conversion between photons [105], relevant for quantum information processing [106] and molecular spectroscopy [107].
Despite the potential for achieving high quantum optomechanical cooperativity in single-molecule host-guest systems, the inherently short picosecond-scale coherence times of molecular vibrations pose significant challenges for practical applications. These short lifetimes limit the ability to perform coherent manipulations and necessitate the use of high-power laser driving [108], ultrafast excitation and detection schemes or deep-subwavelength plasmonic nanocavities [105, 109]. Therefore, extending vibrational lifetimes into the nanosecond regime represents a critical step toward enabling more robust applications.
In traditional cavity optomechanics, similar challenges for acoustic phonons have been addressed using techniques such as impedance mismatch engineering and phononic bandgap structuring, which decouple specific vibrational modes from their environment and thereby prolong coherence. Analogously, in molecular systems, promising routes for achieving efficient optomechanical interactions have been proposed as discussed in Section III.1, even in the presence of modest electron-phonon coupling strengths compared to electron–vibron interactions. However, extending such vibrational control to optical phonons and localized guest-molecule vibrations presents a more formidable challenge due to anharmonic vibrational interactions (see relevant discussions in Section II.2). Potential strategies to tackle this problem include the use of host materials with a sparse vibrational density of states, such as crystalline benzene, the design of molecular systems inherently decoupled from their vibrational surroundings [110], or the exploitation of topological degrees of freedom in molecular systems [111].
Moreover, given their strong electron-vibron coupling and minimal thermal population even at room temperature, engineering long-lived optical phonons and localized guest-molecule vibrational states could enable interesting quantum optomechanical applications at elevated temperatures [112]. Realizing operation above millikelvin temperatures, however, requires not only the suppression of vibrational decay but also the minimization of vibrational dephasing. While negligible dephasing has been demonstrated for certain optical phonon modes in naphthalene molecular crystals up to $60$ K [113], similar behavior remains to be explored in single-molecule host-guest systems. This property may offer an advantage over traditional cavity optomechanical setups, which are prone to dephasing even at millikelvin temperatures due to coupling to two level-system defects [95]. As a result, the molecular approach offers the potential to explore previously inaccessible regimes in optomechanical and nonequilibrium condensed-matter systems.
@@
Figure 6: (a) Achieving optical manipulation of the singlet-triplet transition (red arrow) will provide a “switch” to control the electron spin, allowing it to be turned “on” and “off” on demand. This capability combined with photon scattering via the |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ to |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ ZPL transition (blue arrows) can facilitate optical interface of the electron spin. (b) With an optical cavity resonantly coupled to the ZPL transition, an incoming photon either traverses the cavity when the electron is in the triplet state, or reflects from the front mirror when the electron is in the singlet state.
@@ III.3 Spin-photon interface
The spin states are long-lived and well-isolated from their environment, making them attractive candidates as quantum memory units. Additionally, the excellent properties of molecular ZPL emission offer the possibility to coherently interface spins with propagating photons. These two features combined make it possible to realize a molecular spin-photon interface, which can serve as a key component for quantum networks [114].
To realize an efficient spin-photon interface, it is necessary not only to have optically addressable spin states but also to achieve a coherent coupling of the emitter with a single optical mode. For solid-state emitters, the efficiency of such coherent interaction is constrained by the emitter’s internal quantum efficiency, the branching ratio of ZPL emission, and the optical dephasing rate. Typical PAH molecules used in quantum optical experiments have branching ratio of $\sim$30and low intersystem crossing rate ($<10^{-5}$), when a suitable host matrix is chosen. For the near-infrared-emitting dibenzoterrylene molecule, quantum efficiency larger than 50% has been recently measured [115]. While for molecules emitting in the visible range, higher quantum efficiencies are expected, following the energy-gap law. This combination of deficiencies can be mitigated through resonant coupling to an optical cavity. Using a Fabry-Perot cavity with a $Q$-factor of $2.3\times 10^{5}$ and mode volume of 4.4 $\lambda^{3}$, recent experiments have achieved strong cavity-molecule coupling with cooperativity of 45 [116, 13]. In this regime, emitter-cavity coupling can be combined with spin control to realize a spin-photon interface using deterministic interaction schemes [117].
@@
Figure 6(a) illustrates a possible scheme of a molecular spin-photon interface. To entangle the electron spin with photons, one needs to prepare the molecule in a superposition of |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ and a sublevel of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$. This can in principle be achieved using a narrow linewidth laser to coherently drive the transition between these two states. The spin-photon interaction takes place as follows. As illustrated in Fig. 6(b), the |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ to |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ transition is strongly coupled to an optical microcavity and serves as an optical interface. Depending on the state of the molecule, a photon with its frequency matching the undressed cavity frequency will either reflect off from the cavity, due to the vacuum Rabi splitting [13] when the molecule is in |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, or transmit through the cavity when the molecule is in |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, considering a symmetric mirror configuration. Thus the internal state of the molecule becomes entangled with the spatial mode of the photon.
@@
A critical step in this scheme is the optical control of singlet-triplet transition, as shown by the red arrow in Fig. 6(a), which has not yet been demonstrated on a single PAH molecule. Such transitions have been studied in experiments with cold alkaline earth atoms, e.g., strontium, and utilized for optical clocks transitions [118]. In recent years, singlet-triplet transitions have been measured on diatomic molecules in the gas phase [119, 120]. Compared to the gas-phase experiments, molecules embedded in solids bring along several additional challenges. First, PAH molecules have planar or near-planar structures and do not contain heavy atoms, leading to weak spin-orbit coupling and thus long triplet-state lifetimes. Second, the exact frequencies of these transitions are difficult to predict, and the matrix introduces shifts to the transition frequency compared to that of a molecule in vacuum. Third, the coherent manipulation of singlet-triplet transition needs to surpass deficiencies led by non-radiative decay and pure dephasing resulted from coupling to phonons. Although triplet-state lifetime can be routinely measured, determining other key parameters such as the phosphorescence quantum yield, branching ratio and pure dephasing rate has been rarely reported, apart from an earlier work on perylene in anthracene [121]. A considerable amount of preparatory work in triplet spectroscopy is necessary to unlock the outlined possibilities. In addition, considerations need to be taken in identifying molecular candidates for such investigations. We will discuss possible molecular candidates in Section III.5.
@@ III.4 Molecular spin register
Apart from nuclear spin qubits, using nuclei with larger spin quantum numbers can extend the system to a qudit [^2] A quantum system with a $d$ number of levels. register. Nuclear spins with higher multiplicity offer an interesting route for quantum information processing beyond qubit-based approaches. They are predicted to bring along reduced numbers of physical quantum units and gate operations for accomplishing a certain computing task [129]. A parallel line of research on molecular qudits is being pursued in the context of molecular nanomagnets [129]. On the PAH platform, it is readily feasible to incorporate a defined number of spin-1 nuclei, e.g., 14N, through bay functionalization with imide groups [130]. Along this line, perylenediimide and terrylenediimide derivatives have already been studied in single-molecule optical spectroscopy [131, 132].
Overall, the molecular framework provides the capability to integrate electron spin, nuclear spins, and optical interfaces within a single platform. Compared to well-studied platforms such as semiconductor quantum dots, rare-earth ions and diamond color centers, the proposed system based on PAH molecules operates in different regimes in terms of system dimensions and hyperfine coupling strengths. In quantum dot systems, the central spin is surrounded by a large number ($\sim 10^{5}$) of nuclear spins of the atoms constituting the dot with hyperfine interaction strengths stronger than the linewidth of the electron spin resonance [133]. In this regime, which may be termed the strongly-coupled dense-spin limit, the constituent nuclear spins are indistinguishable to the central spin. The control over individual nuclear spin states is out of reach but collective states of the ensemble can be probed and addressed [134]. Recent advances in rare-earth ion systems, operated in a comparable parameter regime, have demonstrated that a subset of lattice spins can be isolated from the dense ensemble through the frozen-core effect, enabling quantum state preparation within an isolated subset [135]. In contrast, color centers in diamond, such as the nitrogen-vacancy center, provide an implementation of central-spin systems in the dilute-spin limit. The color center can have a few strongly coupled nuclear spins from the nitrogen nucleus forming the center and ${}^{13}\text{C}$ atoms on nearby lattice sites, with hyperfine coupling strengths ranging from 300 kHz to 130 MHz [136]. In addition, weakly coupled nuclear spins, i.e., spins with hyperfine couplings narrower than the electron spin resonance linewidth can also be detected through dynamic decoupling sequences. This enables selective addressing of individual nuclear spins, due to their smaller number and different hyperfine coupling strengths. While it offers advantages in coherence time and controllability, limitations arise from the fact that the majority of addressable nuclear spins are weakly coupled and randomly dispersed within the diamond lattice.
@@
In a PAH molecule, the hyperfine interaction strength of proton and ${}^{13}\text{C}$ spins significantly exceeds the electron spin transition linewidth of $\sim 120\,\text{kHz}$ [77], putting them in the strongly-coupled regime. The molecular framework further allows the spin composition to be varied through chemical modifications, such as perdeuteration. In addition to the dopant molecules, the solid-state environments for PAH molecules, which are normally crystals of hydrocarbon molecules of a smaller size, can also be configured. For example, the transition of a dense-spin to a dilute-spin environment can be achieved through deuteration of the matrix molecules [76], allowing tuning of the coupling strength between the spin register and its environment. In addition to the tailorability by chemical modification, a particularly appealing aspect of this system is the metastable triplet state. Once optical driving of such a transition is realized, it enables a deterministic control to switch “on” and “off” the electron spin [see the red arrow in Fig. 6(a)]. This feature is particularly compelling, given that the electron spin acts as a significant source of decoherence for the nuclear spins [136]. By enabling a deterministic on-off switch of the hyperfine interaction, this system can address the trade-off between gate time and susceptibility to decoherence, further enhancing the utility of the molecular platform.
@@
III.5 Uncovering the Chemical Space
@@
We provide a glimpse into the chemical space of molecules in Fig. 7, where $26$ centrosymmetric PAH molecules containing up to $26$ carbon atoms are identified from the NIST database [137]. Their singlet-singlet (|$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$) and singlet-triplet (|$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$) vertical transition energies are calculated using ab initio methods at the level of coupled-cluster theory with single and double excitations (CCSD) [145] and plotted versus each other. Note that these energies do not include solvation effects or zero-point energy correction stemming from the vibrational degrees of freedom. The results show an approximately linear scaling behavior between the singlet-singlet and singlet-triplet transition energies (see the dashed black line in Fig. 7), spanning a large range in the visible and near-infrared spectrum. This molecular space has not been studied at the single-molecule level except for perylene and pentacene. We discuss the potential of such a diverse chemical space with two possible applications.
@@
(i) The large span of |$\ket{S_{0};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}-\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ transition energies of these molecules offers possibilities to engineer single photon sources with on-demand emission frequencies across the visible spectrum. Although the identified molecules do not cover the entire range, the wavelength gaps among them could be bridged using several strategies. First, chemical modifications, such as the adaptation of imide or anhydride groups, and bay-position functionalization can change the ZPL emission wavelength over an extended range [131, 132]. Second, selecting different host matrices or different insertion sites in a host can affect the emission wavelength by up to $\sim$20 nm [45]. Third, within a given host matrix, the center of inhomogeneous broadening can be tuned with the thickness of the host crystal, providing an accessible frequency range of several nanometers [149]. Fourth, the inhomogeneous broadening of ZPL frequencies can span over $\sim$1 nm, allowing spectral selection of molecules within this frequency range. Finally, fine adjusting of the ZPL emission frequency can be achieved through the Stark effect, with a tunning range of $\sim$500 GHz [16]. Together, these strategies constitute a toolbox for engineering molecular emitters with emission wavelengths tailorable on demand.
@@
(ii) The molecules explored in Fig. 7 offer a valuable chemical space for identifying molecular candidates suitable for realizing spin-photon interfaces, where the optical addressing of the singlet-triplet transition is essential but remains an ongoing challenge [45]. Molecules suitable for addressing singlet-triplet transitions should meet two conditions regarding their transition energy: First, molecules should have a high |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ energy, as dictated by the energy-gap law, meaning that a lower energy of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ is associated with a larger non-radiative ISC rate [81]. Second, the energy of |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ should not exceed visible frequencies significantly, as this would complicate the identification of suitable host materials and necessitate the use of widely tunable laser sources in the ultraviolet, which is technically demanding. Since the energy of |$\ket{T_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$ is typically lower than that of |$\ket{S_{1};\bm{\nu}_{v}=\bm{0},\bm{\nu}_{p}=\bm{0}}$, it is crucial to identify molecules that can balance these energy levels to facilitate efficient spin-photon interactions. These two criteria constrain the selection of suitable guest molecules to the subspace illustrated by the blue-shaded area in Fig. 7.
These two examples are just a glimpse of the immense possibilities offered by chemical space. In the future, significantly larger regions of this space may be explored through machine learning approaches in both theoretical frameworks [150] and experimental workflows [151], enabling the identification of new host-guest combinations with specific properties such as ZPL frequencies at telecom wavelengths, long vibrational lifetimes, strong electron–vibration coupling, extended spin coherence times or high quantum yield emissions from molecular transitions. Organometallic complexes, for example, have already demonstrated the potential to achieve ZPL frequencies at telecom wavelengths [152], showcasing the capabilities that could be realized as more chemical space is explored. In addition, inverse molecular design techniques [148] can be employed to design new molecules with tailored photonic, vibronic, and spin characteristics, harnessing the full potential of chemical synthesis. These approaches can be integrated with advanced synthetic methodologies, such as click chemistry [153] and iterative molecular assembly [154], to enable the realization of custom-designed molecular systems. Ultimately, leveraging generative models and autonomous laboratories could push the boundaries of this chemical space exploration [151, 155]. These strategies hold the promise to transcend traditional design paradigms and significantly accelerate the development of molecules for quantum technologies.